OpenStudy (anonymous):

The set of points (-2, 1), (0, 3), (2, 5) is collinear.

OpenStudy (anonymous):

and?

OpenStudy (anonymous):

is it true or false?

OpenStudy (anonymous):

if you increase the x coordinate, does the y coordinate go up or down in the same ratio for all the points?

OpenStudy (anonymous):

Let A(-2,1), B(0,3) and C(2,5). If the set {A,B,C} is co-linear then : \[\Large \overrightarrow{AB}=k\overrightarrow{AC}~~~~~~\text {where k is real} \]

OpenStudy (anonymous):

So : \[\Large\begin{pmatrix}0-(-2)\\3-1\end{pmatrix}=k\begin{pmatrix}2-(-2)\\5-1\end{pmatrix}\] so : \[\Large \begin{cases}4k=2\\4k=2\end{cases}\] so : \[\Large k=2\] So the set is collinear